OpenStudy (anonymous):

double check my work? A store is having a sale on chocolate chips and walnuts. For 6 pounds of chocolate chips and 2 pounds of walnuts, the total cost is $14 . For 3 pounds of chocolate chips 5 and pounds of walnuts, the total cost is $20. Find the cost for each pound of chocolate chips and each pound of walnuts. I made the equation 6c + 2w = 14, solved for both variables and got c = 2 and w = 1. However, when I substitute those variables into the second equation (3c + 5w = 20), the variables do not work.

OpenStudy (anonymous):

you started out right at the start but you need to solve the 2 equations simultaneously

OpenStudy (anonymous):

should i combine like terms from both equations?

OpenStudy (anonymous):

start with the first equation, you need to either find c or w. So just say we will find c...

OpenStudy (anonymous):

6c + 2w = 14 so 6c = 14 - 2w

OpenStudy (anonymous):

c = 3?

OpenStudy (anonymous):

nevermind, I solved mine a different way. But the equation you wrote would be c = 2?

OpenStudy (anonymous):

actually, find W, its easier and less confusing 4 u

OpenStudy (anonymous):

6c + 2w = 14 so 2w = 14 - 6c

OpenStudy (anonymous):

w = 7 - 3c

OpenStudy (anonymous):

can u see how i got W?

OpenStudy (anonymous):

the you have to sub W= 7-3c into the second equation

OpenStudy (anonymous):

yes, I understand

OpenStudy (anonymous):

3c + 5(7 - 3c) = 20?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

now expand it and find c

OpenStudy (anonymous):

3c + 5(7 - 3c) = 20 3c + 35 - 15c = 20 35 - 12c = 20 -35 -35 -12c/-12 = -15/-12 I think I did something wrong. Is it okay if i got a decimal? 1.25

OpenStudy (anonymous):

yeah that looks ok, now sub c = 1.25 into one of the original formulas to find w